# JavisDiT [Uncaptioned image]: Joint Audio-Video Diffusion Transformer with Hierarchical Spatio-Temporal Prior Synchronization

### v1

@@ D.4.2 Verification on Metric Quality
    The second 500 negative pairs come from separate generations. In particular, we randomly split the 1,000 positive samples into two parts, and select the first 500 samples to prompt AudioLDM2 [27] to generate audios conditioned solely on text captions, without accessing the corresponding videos. This naturally leads to 500 asynchronous video-audio pairs.
@@
Then, we compare our proposed JavisScore with the previously-used AV-Align [60] metric on the 3,000 evaluation samples, and the result is displayed in Tab. A4. We calculate the AV-Align scores and our JavisScore on all the 3000 samples, and compute the AUROC value for the binary classification (ideally, a VA-synchrony metric should produce higher scores on positive video-audio pairs), which shows JavisScore achieves approximately 0.13 higher AUROC than AV-Align. Then, as each positive sample is related to 2 negative video-audio pairs, we can calculate the prediction accuracy of $1000\times 2=2000$ paired positive-negative samples. It is viewed as correct when a metric assigns a higher score on the positive sample than the corresponding negative video-audio pair. According to Tab. A4, our metric significantly surpasses AV-Align by 23% accuracy, demonstrating the efficacy of video-audio synchrony measurement.
Table A4: Comparison of VA-Synchrony Metrics
@@
Notably, AV-Align nearly performs as a random-guess (with a near 0.5 AUROC and Accuracy). This is because AV-Align simply uses optical flow to capture the video dynamics, and match with audio dynamics estimated by onset detection results. If there is a concurrent “pulse” at the video and audio timelines, AV-Align will produce a high synchrony score. However, in real-world scenarios, the video optical flow and audio onset detection cannot precisely capture the start and end of the visual-audio event. As exemplified in Fig. A5, when a person in the video speaks, the movements of their mouth are subtle and often too minimal to be effectively captured by optical flow. Moreover, in environments with strong background noise (e.g., TV programs), the onset of human talking also becomes difficult to detect. The scores generated by AV-Align are unreliable in such complex scenarios. In contrast, our proposed JavisScore leverages the high-dimensional semantic space of ImageBind [11] to compute synchronization at the second-level, robustly distinguishing between synchronized and unsynchronized cases. It is worth noting, however, that our metric does not achieve 100% accuracy. Developing a more precise evaluation metric remains a significant challenge in the JAVG domain, and we hope this work will inspire further advancements in this critical area.
@@ D.5 In-depth Analysis of Evaluation Results
Table A5: Evaluation on audio generation. After sufficient training iterations at stage 1, our JavisDiT presents moderate audio generation performance on in-domain test set AudioCaps [20] and comparable quality on out-of-domain test set JavisBench-mini.
Method AudioCaps (InD) JavisBench-mini (OoD)
[…]
AudioLDM2 [27] 2.01 0.205 0.487 4.62 0.198 0.428
JavisDiT-audio-ep13 5.88 0.145 0.319 5.11 0.165 0.401
JavisDiT-audio-ep55 5.19 0.164 0.356 4.68 0.194 0.432
@@ Appendix E Additional Experiments
E.1 Evaluation on Audio Generation Quality
@@
  • AudioCaps test set [20]: Filtered by AudioLDM2 [27], this dataset consists of 964 samples, each 8–10 seconds long. Models are required to generate 10-second audio clips.
@@
On these datasets, we first analyze the performance progression of our model across different training stages and then compare it against the current state-of-the-art (SOTA) model, AudioLDM2 [27].
As shown in Tab. A5, with increased training iterations (from epoch 13 to epoch 55), our model demonstrates improvements in both audio generation quality (e.g., FAD decreases from 5.88 to 5.19 on AudioCaps) and text consistency (e.g., TA-IB increases from 0.165 to 0.194, while CLAPScore improves from 0.401 to 0.432 on JavisBench-mini). Furthermore, our model achieves comparable audio generation performance to AudioLDM2 on JavisBench-mini, further validating the effectiveness of our DiT-based architecture.
Notably, our model performs worse than AudioLDM2 on AudioCaps, which can be attributed to two primary factors: (1) Training strategy: AudioLDM2 is explicitly trained for 10-second audio generation, whereas our model supports variable-length audio generation. Consequently, our model currently lacks sufficient training data and iterative steps specifically for 10-second audio generation. (2) Potential data leakage: The training data of AudioLDM2 includes the training set of AudioCaps, leading to a smaller in-domain test gap, which increases the risk of overfitting. In contrast, JavisBench is collected from more diverse real-world YouTube audio sources, and its domain-specific soundscapes are not exposed to either AudioLDM2 or our model during training. Thus, performance on the out-of-domain JavisBench-mini provides a more realistic reflection of a model’s usability in real-world scenarios.
@@
E.2 Investigation of ST-Prior’s Extraction and Representation
@@
$\surd$ $\times$ $\times$ $\times$ 0.190 0.799 0.167 0.097 0.149
$\surd$ $\surd$ $\times$ $\times$ 0.193 0.799 0.170 0.102 0.151
$\surd$ $\surd$ $\surd$ $\times$ 0.196 0.800 0.174 0.096 0.156
$\surd$ $\surd$ $\times$ $\surd$ 0.202 0.800 0.179 0.107 0.159
$\surd$ $\surd$ $\surd$ $\surd$ 0.211 0.801 0.190 0.122 0.169
Next, using the 1:1 ratio, we explored different configurations of prior numbers and dimensions. The results in Fig. A8 indicate that increasing the prior number and dimension consistently improves performance, demonstrating the scalability of our approach. In this work, we chose the “n32x32+d128” configuration as it offers a good trade-off between performance and training cost, considering the diminishing returns of further scaling.
E.3 Training Losses for ST-Prior’s Estimation
@@
On the other hand, $\mathcal{L}_{vad}$ considerably enhances the synchrony of generated video-audio pairs (e.g., 0.156 vs. 0.151 on JavisScore), thanks to its ability to maximize the divergence between positive and negative video-audio samples themselves. However, $\mathcal{L}_{reg}$ offers even greater benefits due to its smoother regularization, facilitating the convergence of the text prior to the positive embeddings.
Moreover, the combination of all loss functions achieves the best performance (e.g., 0.169 of JavisScore), as they collaboratively address the same goal from different perspectives: embedding video-audio synchrony into the text prior and ensuring it captures the semantic meaning of synchronization. This comprehensively validates both our motivation and the effectiveness of our methodology.
E.4 More Generation Examples
@@
Appendix F Extension to X-conditional Generation
F.1 Masked Training for Joint Video-Audio Generation
@@
Built on diffusion models with transformers (DiT), our JavisDiT can be easily extended to support various conditional video-audio generation tasks, as shown in Fig. A10. Inspired by UL2 [49] and OpenSora [65], we propose a dynamic masking strategy to support video and audio conditioning, from the basic (I) text-to-audio-video (t2av) generation to (II) audio-to-video (a2v) generation, (III) video-to-audio (v2a) generation, (IV) audio-image-to-video (ai2v) generation, (V) image-to-audio-video (i2av) generation, and (VI) audio-video-extension (av_ext) generation. Note that the text condition still works in all kinds of conditional generations, which provides both coarse-grained global semantic embeddings and fine-grained spatio-temporal priors on targeted visual-sounding events. In particular, we unmask the specific video and frames to be conditioned on for X-conditional generation. During both model training and inference, unmasked frames will have timestep 0, while others remain the same (t):
@@
We randomly apply one of the last five (from (II) to (IV)) masking strategy during training with a total probability of 30%, and perform the original text-to-audio-video generation task on the remaining 70% samples.
F.2 Application to X-conditional Generation
Fig. A11 illustrates a list of representative x-conditional cases:
  • A2V/V2A: In the audio-to-video (A2V) mode, due to the inherent randomness of the diffusion process, the model generates diverse outputs while maintaining spatial consistency (e.g., a man in blue playing a blues guitar piece) and temporal synchronization (e.g., the sound is produced precisely when the fingers strum the strings). Similarly, in the video-to-audio (V2A) mode, although the generated audio differs from the ground truth, it retains high quality, textual consistency, and precise video-audio synchronization.
  • AI2V/I2AV: When given the first frame of an image (I2V), our model can animate the image, generating videos that resemble the ground truth (GT) more closely while being distinct from those in A2V. In cases where both audio and the initial image are provided (AI2V), the generated video aligns more closely with the ground truth. Conversely, when only an image is given (I2AV), the generated audio resembles the V2A results.
  • AV-Ext: When provided with a short clip (e.g., 2 seconds) of audio and video, our model effectively understands the preceding content and generates semantically and contextually consistent continuations. This capability is particularly beneficial for generating long videos [65, 36]. By employing a stepwise extension strategy, the model mitigates challenges such as memory overflow and difficulties in ensuring continuity when directly generating longer videos (e.g., 1 minute).
Thanks to our proposed dynamic masking strategy, in addition to standard JAVG tasks, our model also supports various modality-conditional generation tasks, opening up new possibilities for a wide range of applications within the research community.
Figure A11: Zero-shot X-conditional generation cases. Our model ensures diversified audio-video generation while accurately generating based on various conditions. (1) For A2V/V2A tasks, our model’s generation differs from ground-truth video/audio but maintains spatial consistency and temporal synchronization (similar finger movements and approximate start/end timestamps). (2) For AI2V/I2AV tasks, we nearly recover the ground-truth video in AI2V while keep similar audio results between I2AV and V2A. (3) For AV-Ext, our model produces consistent continues but differs from both GT-AV and I2AV results.

### v2

@@ D.4.2 Verification on Metric Quality
    The second 500 negative pairs come from separate generations. In particular, we randomly split the 1,000 positive samples into two parts, and select the first 500 samples to prompt AudioLDM2 (Liu et al., 2024b) to generate audios conditioned solely on text captions, without accessing the corresponding videos. This naturally leads to 500 asynchronous video-audio pairs.
@@
Figure 4: Comparison of VA-synchrony metrics. $\dagger$ means we relax the criteria to $Acc=\mathbb{I}\{s^{+}\geq s^{-}\}$.
@@
Then, we compare our proposed JavisScore with the previously-used AV-Align (Yariv et al., 2024) metric on the 3,000 evaluation samples, and the result is displayed in Fig. 4. We calculate the AV-Align scores and our JavisScore on all the 3000 samples, and compute the AUROC value for the binary classification (ideally, a VA-synchrony metric should produce higher scores on positive video-audio pairs), which shows JavisScore achieves approximately 0.13 higher AUROC than AV-Align. Then, as each positive sample is related to 2 negative video-audio pairs, we can calculate the prediction accuracy of $1000\times 2=2000$ paired positive-negative samples. It is viewed as correct when a metric assigns a higher score to the positive sample than to the corresponding negative video-audio pair. According to Fig. 4, our metric significantly surpasses AV-Align by 23% accuracy, demonstrating the efficacy of video-audio synchrony measurement.
@@
Notably, AV-Align nearly performs as a random guess (with a near 0.5 AUROC and Accuracy). This is because AV-Align simply uses optical flow to capture the video dynamics, and match with audio dynamics estimated by onset detection results. If there is a concurrent “pulse” at the video and audio timelines, AV-Align will produce a high synchrony score. However, in real-world scenarios, the video optical flow and audio onset detection cannot precisely capture the start and end of the visual-audio event. As exemplified in Fig. A7, when a person in the video speaks, the movements of their mouth are subtle and often too minimal to be effectively captured by optical flow. Moreover, in environments with strong background noise (e.g., TV programs), the onset of human talking also becomes difficult to detect. The scores generated by AV-Align are unreliable in such complex scenarios. In contrast, our proposed JavisScore leverages the high-dimensional semantic space of ImageBind (Girdhar et al., 2023) to compute synchronization at the second-level, robustly distinguishing between synchronized and unsynchronized cases. It is worth noting, however, that our metric does not achieve 100% accuracy. Developing a more precise evaluation metric remains a significant challenge in the JAVG domain, and we hope this work will inspire further advancements in this critical area.
In addition, SynchFormer (Iashin et al., 2024) employs a 21-category classifier to estimate the temporal shift degree between video and audio, which is then used to compute a synchrony score (lower is better). However, as shown in Fig. 4, its discrete predictions result in poor separation between positive and negative samples, yielding low accuracy and AUROC. Only when we relax the evaluation metric to a $Acc=\mathbb{I}\{s^{+}\geq s^{-}\}$ criterion does its accuracy approach that of JavisScore. Besides, SynchFormer is inherently limited to predicting temporal shifts and is not applicable to the diverse spatiotemporal misalignment scenarios considered in the real world (see Sec. C.2.4). Nevertheless, its approach suggests a promising direction: learning a neural network–based synchrony metric via supervision, which we leave as future work.
@@ Appendix E Additional Experiments
E.1 Full Evaluation Results for Ablation Studies
In the manuscript, we use several proxy metrics ($S_{\text{AVQ}}$; $S_{\text{AVC}}$; $S_{\text{AVS}}$) to simplify the presentation when evaluating the quality, consistency, and synchrony of generated audio–video outputs. In this section, we provide the full results corresponding to Tab. 4 (model-design ablations) and Tab. 4 (ST-Prior ablations) in Tab. A5 and Tab. A6, respectively, to further enhance the clarity and completeness of the paper.
Table A5: Detailed evaluation results of ablation on model design (Tab. 4.)
STDiT HiST-Sypo BiCA AV-Quality Text-Consistency AV-Consistency AV-Synchrony
[…]
$\times$ $\times$ $\times$ 442.0 3.9 10.5 0.268 0.139 0.287 0.344 0.197 0.796 0.147 0.118
$\surd$ $\times$ $\times$ 335.4 3.4 8.1 0.276 0.145 0.309 0.380 0.200 0.799 0.156 0.130
$\surd$ $\surd$ $\times$ 249.7 2.9 7.4 0.275 0.147 0.305 0.375 0.207 0.800 0.184 0.150
$\surd$ $\times$ $\surd$ 269.1 3.1 7.9 0.273 0.144 0.300 0.377 0.200 0.795 0.162 0.133
$\surd$ $\surd$ $\surd$ 241.8 2.9 7.3 0.277 0.146 0.308 0.382 0.209 0.801 0.186 0.153
Table A6: Detailed evaluation results of ablation on ST-Prior design (Tab. 4.)
$\bm{N_{s}}$ $\bm{N_{t}}$ Injection AV-Quality Text-Consistency AV-Consistency AV-Synchrony
[…]
0 0 - 269.1 3.1 7.9 0.273 0.144 0.300 0.377 0.200 0.795 0.162 0.133
1 1 CrossAttn 289.9 3.2 8.1 0.269 0.140 0.293 0.375 0.204 0.797 0.175 0.137
16 16 CrossAttn 257.2 3.0 7.5 0.275 0.144 0.305 0.381 0.209 0.801 0.185 0.151
32 32 CrossAttn 241.8 2.9 7.3 0.277 0.146 0.308 0.382 0.209 0.801 0.186 0.153
32 32 Addition 252.7 3.0 7.4 0.274 0.142 0.303 0.380 0.206 0.798 0.180 0.144
32 32 Modulate 255.0 2.9 7.4 0.277 0.143 0.309 0.382 0.207 0.800 0.180 0.145
Table A7: Evaluation on audio generation. After sufficient training iterations at stage 1, our JavisDiT presents moderate audio generation performance on in-domain test set AudioCaps (Kim et al., 2019) and comparable quality on out-of-domain test set JavisBench-mini.
Method AudioCaps (InD, 10s) AudioCaps (InD, 4s) JavisBench-mini (OoD, 4s)
[…]
AudioLDM2 2.01 0.205 0.487 5.57 0.147 0.326 8.81 0.153 0.360
JavisDiT-audio-ep13 5.88 0.145 0.319 - - - 8.33 0.150 0.368
JavisDiT-audio-ep55 5.19 0.164 0.356 6.23 0.141 0.301 8.11 0.152 0.381
E.2 Evaluation on Audio Generation Quality
@@
  • AudioCaps test set (Kim et al., 2019): Filtered by AudioLDM2 (Liu et al., 2024b), this dataset consists of 964 samples, each 8–10 seconds long. Models are required to generate 10-second and 4-second audio clips.
@@
On these datasets, we first analyze the performance progression of our model across different training stages and then compare it against the current state-of-the-art (SOTA) model, AudioLDM2 (Liu et al., 2024b).
As shown in Tab. A7, with increased training iterations (from epoch 13 to epoch 55), our model demonstrates improvements in both audio generation quality (e.g., FAD decreases from 5.88 to 5.19 on AudioCaps) and text consistency (e.g., TA-IB increases from 0.145 to 0.164, while CLAPScore improves from 0.368 to 0.381 on JavisBench-mini). Furthermore, our model achieves comparable audio generation performance to AudioLDM2 on JavisBench-mini, further validating the effectiveness of our DiT-based architecture.
Notably, our model performs worse than AudioLDM2 on AudioCaps, which can be attributed to two primary factors: (1) Training strategy: AudioLDM2 is explicitly trained for 10-second audio generation, whereas our model supports variable-length audio generation. In particular, although one can forcibly alter AudioLDM2’s generation length during inference by directly modifying the shape of the noisy latent (e.g., changing it to 4 seconds), the model’s denoising process becomes misaligned with its training distribution, leading to a significant drop in generation quality, as shown in Tab. A7. In contrast, since our audio pretraining is performed under variable-length settings, the generation quality at 4 seconds remains much closer to that of AudioLDM2. We believe that collecting more data and training for additional iterations will further improve generation performance across different lengths. (2) Potential overfitting problem: The training data of AudioLDM2 includes the training set of AudioCaps, leading to a smaller in-domain test gap, which increases the risk of overfitting. In contrast, JavisBench is collected from more diverse real-world YouTube audio sources, and its domain-specific soundscapes are not exposed to either AudioLDM2 or our model during training. Thus, performance on the out-of-domain JavisBench-mini provides a more realistic reflection of a model’s usability in real-world scenarios.
@@
E.3 Investigation of ST-Prior’s Extraction and Representation
Next, using the 1:1 ratio, we explored different configurations of prior numbers and dimensions. The results in Fig. A10 indicate that increasing the prior number and dimension consistently improve performance, demonstrating the scalability of our approach. In this work, we chose the “n32x32+d128” configuration as it offers a good trade-off between performance and training cost, considering the diminishing returns of further scaling.
@@ E.3 Investigation of ST-Prior’s Extraction and Representation
$\surd$ $\times$ $\times$ $\times$ 0.190 0.799 0.167 0.097 0.133
$\surd$ $\surd$ $\times$ $\times$ 0.193 0.799 0.170 0.102 0.136
$\surd$ $\surd$ $\surd$ $\times$ 0.196 0.800 0.174 0.096 0.140
$\surd$ $\surd$ $\times$ $\surd$ 0.202 0.800 0.179 0.107 0.149
$\surd$ $\surd$ $\surd$ $\surd$ 0.209 0.801 0.186 0.122 0.153
E.4 Training Losses for ST-Prior’s Estimation
@@
On the other hand, $\mathcal{L}_{vad}$ considerably enhances the synchrony of generated video-audio pairs (e.g., 0.140 vs. 0.136 on JavisScore), thanks to its ability to maximize the divergence between positive and negative video-audio samples themselves. However, $\mathcal{L}_{reg}$ offers even greater benefits due to its smoother regularization, facilitating the convergence of the text prior to the positive embeddings.
Moreover, the combination of all loss functions achieves the best performance (e.g., 0.153 of JavisScore), as they collaboratively address the same goal from different perspectives: embedding video-audio synchrony into the text prior and ensuring it captures the semantic meaning of synchronization. This comprehensively validates both our motivation and the effectiveness of our methodology.
E.5 Additional Visualization of ST-Prior Injection
Figure A11: Attention shifts over sequential audiovisual events during ST-Prior injection.
In Fig. A11, we include a more detailed generation example to illustrate how attention shifts during sequential ST-Prior injection:
  • First, when the orange cat on the left meows, the visual spatial attention concentrates on its head, which corresponds to the region highlighted by the white dashed box in the audio-attention map.
  • Then, when the tabby cat on the right meows (indicated by the red dashed box in the audio spectrogram), the visual attention shifts to the tabby’s head, providing a more direct spatial prior.
  • Finally, when the orange cat meows again, both spatial and temporal attention shift once more, injecting updated prior information into the audio and video branches.
This example more clearly demonstrates how spatiotemporal priors are injected and shifted over sequential events, further validating the effectiveness of our method.
E.6 More Generation Examples
@@
E.7 Discussion: Extension to X-conditional Generation
@@
Built on diffusion models with transformers (DiT), our JavisDiT can be easily extended to support various conditional video-audio generation tasks, as shown in Fig. A13. Inspired by UL2 (Tay et al., 2022) and OpenSora (Zheng et al., 2024), we propose a dynamic masking strategy to support video and audio conditioning, from the basic (I) text-to-audio-video (t2av) generation to (II) audio-to-video (a2v) generation, (III) video-to-audio (v2a) generation, (IV) audio-image-to-video (ai2v) generation, (V) image-to-audio-video (i2av) generation, and (VI) audio-video-extension (av_ext) generation. Note that the text condition still works in all kinds of conditional generations, which provides both coarse-grained global semantic embeddings and fine-grained spatio-temporal priors on targeted visual-sounding events. In particular, we unmask the specific video and frames to be conditioned on for X-conditional generation. During both model training and inference, unmasked frames will have timestep 0, while others remain the same (t):
@@
We view the integration of all audio-video interactive generation tasks in one unified model as our future work.